OpenStudy (anonymous):

Help please!! will give medals A confectionary company did a survey among students. It was found that 80 students liked pastries and 30 of those students liked candies also. There were 15 students who did not like pastries but liked candies. There were 5 students who did not like either pastries or candies. Using a two-way table, predict the number of students who would like candies when 200 students participate in the survey. 45 50 90 95

OpenStudy (anonymous):

@CGGURUMANJUNATH do you mind helping us out?

OpenStudy (anonymous):

um i addedd up all the numbers and got 130 not the extra numbers but i have to subtract sooo

OpenStudy (anonymous):

um i got 65 tht may not be right

OpenStudy (anonymous):

i'll show my work correct me if i'm wrong

OpenStudy (anonymous):

a solution exists!

OpenStudy (anonymous):

wow this is really hard

OpenStudy (anonymous):

@agent0smith can you help us out???

OpenStudy (anonymous):

yeah thats right

OpenStudy (anonymous):

oh the extra numbers!!! this is a choice question/bubble in makes it all the more easier

OpenStudy (anonymous):

try to mark off numbers as we go

OpenStudy (anonymous):

this is an obvious yet not obvious q

OpenStudy (anonymous):

i'll start again but do it differently, 30 of the 80 like candy so it may not be 90/95 but then again 200 people did the survey

OpenStudy (anonymous):

sour wing i need your help

OpenStudy (anonymous):

tell me what you have so far i'm gonna do this the hard way

OpenStudy (anonymous):

i got 90

OpenStudy (anonymous):

ok that makes since

OpenStudy (anonymous):

wait no thats not it!

OpenStudy (anonymous):

because 200-130=70 right? but i'm not saying its 70

OpenStudy (anonymous):

because you should subtract the 5 that don't like candy but to figure that out i need you to tell me how you did the work

OpenStudy (anonymous):

wait i have an idea

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

how about you take that 30 from the 130 and subtract/add that later on you could do it like this:200-100=100-5=95

OpenStudy (anonymous):

but the 30 would have to be added or -ed

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yeah cause tht puts us in a corner

OpenStudy (anonymous):

45/100 students like candies yes? so if there 200 students, 200*45/100=90 students like candies

OpenStudy (anonymous):

yeah he's right!

OpenStudy (anonymous):

i think your right! it makes sense

OpenStudy (anonymous):

wow if only we had done that before........